# The asymptoticity of extremal length in Teichmüller space

### v1

@@ 1. Introduction
The limit is infinite if the measured foliation $F$ intersects $V(q)$, since in this case the measured foliation $F^{\prime}$ can be chosen to coincide with $V(q)$. Moreover, if the measured foliation $F$ is similar to $V(q)$, that is, if $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$ is a coefficient, the right side of the formula can be simplified.
@@
Let $\mathcal{R}_{q,X}(t)$ be the Teichmüller ray determined by the unit-norm quadratic differential $q$, and let $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ be the vertical foliation. If a measured foliation $F\in\mathcal{MF}$ can be written in the form $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$ is a coefficient, then
@@
For bounded Teichmüller rays, the remaining questions concern the limiting Teichmüller distances of pairs of Teichmüller rays and the conditions under which they are asymptotic. In [13], Masur considered the case of uniquely ergodic foliations, proving that two bounded Teichmüller rays with uniquely ergodic vertical foliations are asymptotic if their foliations contain no simple closed curves formed by saddle connections. Amano [3] derived a formula for the limiting Teichmüller distance of Jenkins-Strebel rays, showing that the limiting distance of bounded Jenkins-Strebel rays depends on both the moduli of cylindrical components of initial surfaces and the Teichmüller distance between their endpoints in the augmented Teichmüller spaces. In [7], we extended this formula to Teichmüller rays along which the surfaces can be decomposed into cylinders and minimal domains whose transverse measures are uniquely ergodic, with the decompsition determined by the leaves of the vertical foliations. By means of Kerckhoff’s formula, we determine the limiting distance for arbitrary pairs of Teichmüller rays through the asymptotic behavior of extremal lengths, thereby establishing an explicit expression for the limit for any such pair.
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by the unit norm quadratic differentials $q$ and $q^{\prime}$ respectively. If the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous, then the minimum of the limiting Teichmüller distance is $\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})$, where $\delta$ is the detour metric and $\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}$ are the Busemann points on the horofunction boundary of $\mathcal{T}(S)$ associated to $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ respectively.
@@ 2.2. Measured foliations
holds for any $\alpha\in\mathcal{S}$. We denote by $F$ the equivalent class containing $(\mathcal{F},\mu)$. Therefore, the function $i(F,\cdot)$ is an element of $\mathbb{R}^{\mathcal{S}}_{\geq 0}$, where $\mathbb{R}^{\mathcal{S}}_{\geq 0}$ is the space of non-negative functions on $\mathcal{S}$ topologized by pointwise convergence. We denote by $\mathcal{MF}$ the space of equivalent classes of measured foliations on $S$ equipped with the weak topology induced by the intersection functions in $\mathbb{R}^{\mathcal{S}}_{\geq 0}$. The intersection function can extend continuously to $\mathcal{MF}\times\mathcal{MF}$ (cf. [4] and [18]).
@@ 2.5. Teichmüller rays
Consider the quasiconformal mappings between two Riemann surfaces $X,Y\in\mathcal{T}(S)$. Teichmüller’s theorem states that there exists a unique quasiconformal mapping that realizes the Teichmüller distance $d_{\mathcal{T}}(X,Y)$, known as the Teichmüller mapping. For the Teichmüller mapping $f:X\to Y$, its Beltrami coefficient $\mu_{f}$ has the form $\mu_{f}=\frac{K(f)-1}{K(f)+1}\frac{\bar{q}}{|q|}$, where $q$ is a unit norm holomorphic quadratic differential on $X$.
Let $f_{q,t}$ be the Teichmüller mapping from $X$ to $X_{t}$ determined by the unit norm holomorphic quadratic differential $q$ on $X$. There is a unit norm holomorphic quadratic differential $q_{t}$ on $X_{t}$, called the terminal differential. Then, in the canonical coordinate $z=x+iy$ of $q$ and the canonical coordinate of $q_{t}$, the Teichmüller mapping $f_{q,t}$ is given by
@@
The Teichmüller geodesic $\mathcal{G}_{q,X}(t)$ through $X$ induced by a unit norm holomorphic quadratic differential $q$ can be defined by
@@ 3. The asymptotic behavior of extremal length
Lemma 3.1 ([21]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
@@
Theorem 3.2 ([21]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit norm holomorphic quadratic differential $q$ on $X$, and $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
@@
If the measured foliation $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$, then, by the Cauchy-Schwarz inequality, there is
@@
for each indecomposable component $G_{j}$. This $F^{\prime}$ satisfying the condition can be obtained by appropriately adjusting the transverse measures of $H(q)$ in the intersection set with $G_{j}$, scaled by the factor $c_{j}/a_{j}$. Then,
@@ 4. The asymptotic behavior of pairs of Teichmüller rays
If two measured foliations $F,F^{\prime}\in\mathcal{MF}$ are absolutely continuous, then it is evident that
@@
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray determined by a unit norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. We define a function $\mathcal{E}_{q}:\mathcal{MF}\to\mathbb{R}_{\geq 0}$ associated with the Teichmüller ray, which is given by
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
Lemma 4.2 ([21]).
Let $q$ and $q^{\prime}$ be two unit norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}>0$. If the vertical foliation of $q^{\prime}$ can be expressed as $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$, where $b_{j}\geq 0$, then
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
Let $q$ and $q^{\prime}$ be two unit norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
where $\mathcal{MF}^{\prime}_{0}(V(q))=\{F\in\mathcal{MF}\mid F=\sum_{j=1}^{n}c_{j}G_{j},\;c_{j}\geq 0\}$.
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then

### v2

@@ 1. Introduction
The limit is infinite if the measured foliation $F$ intersects $V(q)$, since in this case the measured foliation $F^{\prime}$ can be chosen to coincide with $V(q)$. Moreover, if the measured foliation $F$ is similar to $V(q)$, that is, if $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}>0$ is a coefficient, the right side of the formula can be simplified.
@@
Let $\mathcal{R}_{q,X}(t)$ be the Teichmüller ray determined by the unit-norm quadratic differential $q$, and let $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ be the vertical foliation. If a measured foliation $F\in\mathcal{MF}$ can be written in the form $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}>0$ is a coefficient, then
@@
For bounded Teichmüller rays, the remaining questions concern the limiting Teichmüller distances of pairs of Teichmüller rays and the conditions under which they are asymptotic. In [13], Masur considered the case of uniquely ergodic foliations, proving that two bounded Teichmüller rays with uniquely ergodic vertical foliations are asymptotic if their foliations contain no simple closed curves formed by saddle connections. Amano [3] derived a formula for the limiting Teichmüller distance of Jenkins-Strebel rays, showing that the limiting distance of bounded Jenkins-Strebel rays depends on both the moduli of cylindrical components of initial surfaces and the Teichmüller distance between their endpoints in the augmented Teichmüller spaces. In [7], we extended this formula to Teichmüller rays along which the surfaces can be decomposed into cylinders and minimal domains whose transverse measures are uniquely ergodic, with the decomposition determined by the leaves of the vertical foliations.
In [7], we adopted an approach based on constructing quasiconformal mappings between surfaces along two bounded Teichmüller rays, combining techniques due to Amano and Masur. However, this method does not readily extend to the case of Teichmüller rays whose associated vertical foliations contain non-uniquely ergodic minimal components. In this paper, we instead study the limiting Teichmüller distances via Kerckhoff’s formula. This approach allows us to treat the remaining case of non-uniquely ergodic foliations and avoids the technically demanding construction of quasiconformal mappings.
Since the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous for the two bounded Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$, each measured foliation $F\in\mathcal{MF}$ either intersects both $V(q)$ and $V(q^{\prime})$ or intersects neither. Let $\mathcal{MF}_{1}(V(q))$ denote the collection of measured foliations that intersect $V(q)$, and let $\mathcal{MF}_{0}(V(q))$ denote the collection of measured foliations that have zero intersection with $V(q)$. We investigate the ratios of extremal lengths of measured foliations along the two bounded Teichmüller rays and obtain two equalities for the limits of the supremum of these ratios, taken respectively over $\mathcal{MF}_{1}(V(q))$ and $\mathcal{MF}_{0}(V(q))$ (see Theorem 4.3 and Theorem 4.5).
Combining these results with Kerckhoff’s formula, we obtain the limiting distance for arbitrary pairs of Teichmüller rays through the asymptotic behavior of extremal lengths, thereby establishing an explicit expression for the limit for any such pair.
@@
Remark.
In [7], We defined the limiting surfaces for Teichmüller rays and the Teichmüller distance between these limiting surfaces. A comparison between the main result of [7] and Theorem 1.3 shows that, when considering the limit distance for two bounded Teichmüller rays, it is unnecessary to analyze the distance between their limiting surfaces. Indeed, since the intersection number of a simple closed curve with $V(q)$ is preserved along the Teichmüller ray $\mathcal{R}_{q,X}(t)$, each simple closed curve in $\mathcal{MF}_{0}(V(q))$ corresponds naturally to a simple closed curve on the limiting surface of $\mathcal{R}_{q,X}(t)$, which may be homotopic to a puncture. Then, by combining Kerckhoff’s formula with Theorem 4.5, Theorem 1.3 implies that the distance between limiting surfaces of two bounded Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ is bounded by $\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}$.
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by the unit-norm quadratic differentials $q$ and $q^{\prime}$ respectively. If the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous, then the minimum of the limiting Teichmüller distance is $\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})$, where $\delta$ is the detour metric and $\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}$ are the Busemann points on the horofunction boundary of $\mathcal{T}(S)$ associated to $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ respectively.
@@ 2.2. Measured foliations
holds for any $\alpha\in\mathcal{S}$. Equivalently, they are Whitehead equivalence, meaning that one can be transformed into the other by a finite sequence of homeomorphisms of the surface isotopic to the identity, which preserve the transverse measures, together with Whitehead moves, where a Whitehead move is a deformation that collapses a leaf joining two singularities to a point, which may be a puncture, as well as the inverse move of such a collapse.
We denote by $F$ the equivalent class containing $(\mathcal{F},\mu)$. Therefore, the function $i(F,\cdot)$ is an element of $\mathbb{R}^{\mathcal{S}}_{\geq 0}$, where $\mathbb{R}^{\mathcal{S}}_{\geq 0}$ is the space of non-negative functions on $\mathcal{S}$ topologized by pointwise convergence. We denote by $\mathcal{MF}$ the space of equivalent classes of measured foliations on $S$ equipped with the weak topology induced by the intersection functions in $\mathbb{R}^{\mathcal{S}}_{\geq 0}$. The intersection function can extend continuously to $\mathcal{MF}\times\mathcal{MF}$ (cf. [4] and [18]).
@@ 2.5. Teichmüller rays
Consider the quasiconformal mappings between two Riemann surfaces $X,Y\in\mathcal{T}(S)$. Teichmüller’s theorem states that there exists a unique quasiconformal mapping that realizes the Teichmüller distance $d_{\mathcal{T}}(X,Y)$, known as the Teichmüller mapping. For the Teichmüller mapping $f:X\to Y$, its Beltrami coefficient $\mu_{f}$ has the form $\mu_{f}=\frac{K(f)-1}{K(f)+1}\frac{\bar{q}}{|q|}$, where $q$ is a unit-norm holomorphic quadratic differential on $X$.
Let $f_{q,t}$ be the Teichmüller mapping from $X$ to $X_{t}$ determined by the unit-norm holomorphic quadratic differential $q$ on $X$. There is a unit-norm holomorphic quadratic differential $q_{t}$ on $X_{t}$, called the terminal differential. Then, in the canonical coordinate $z=x+iy$ of $q$ and the canonical coordinate of $q_{t}$, the Teichmüller mapping $f_{q,t}$ is given by
@@
The Teichmüller geodesic $\mathcal{G}_{q,X}(t)$ through $X$ induced by a unit-norm holomorphic quadratic differential $q$ can be defined by
@@ 3. The asymptotic behavior of extremal length
Lemma 3.1 ([21, Lemma 3]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit-norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
@@
Theorem 3.2 ([21, Theorem 1]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit-norm holomorphic quadratic differential $q$ on $X$, and $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
@@
If the measured foliation $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}>0$, then, by the Cauchy-Schwarz inequality, there is
@@
for each indecomposable component $G_{j}$. This $F^{\prime}$ satisfying the condition can be obtained by appropriately adjusting the horizontal foliation $H(q)$ in the intersection set with $G_{j}$. Let $\Omega$ be a minimal domain of $V(q)$, and fix a horizontal arc $\tau$ in $\Omega$. Any leaf starting on $\tau$ either terminates at a singularity of $V(q)$ or returns to $\tau$. Let $\mu_{j}$, $j\in J_{\Omega}\subseteq\{1,\cdots,n\}$, denote the ergodic transverse measure of $G_{j}$ supported on $\Omega$. Since the transverse measures $\{\mu_{j}\}_{j\in J_{\Omega}}$ are mutually singular, the horizontal arc $\tau$ admits a decomposition
$\tau=\bigcup_{j\in J_{\Omega}}\tau_{j},$
where $\{\tau_{j}\}_{j\in J_{\Omega}}$ are pairwise disjoint Borel subsets such that $\mu_{j}(\tau_{k})=\mu_{j}(\tau)$ when $j=k$ and $\mu_{j}(\tau_{k})=0$ otherwise.
Consider the lengths of leaves in $\Omega$ with respect to the flat metric induced by $q$. For each $j\in J_{\Omega}$, we rescale the lengths of leaves crossing $\tau_{j}$ by the factor $c_{j}/a_{j}$. Since each leaf in $\Omega$ has infinite length, this rescaling preserves the underlying foliation structure of $V(q)$ on $\Omega$. Similarly, for each cylindrical component $G_{j}$ of $V(q)$, we rescale the lengths of the closed leaves in $G_{j}$ by the factor $c_{j}/a_{j}$. Then, the lengths of saddle connections on the boundaries of cylinders and minimal domains are also rescaled by the corresponding factors.
Since a saddle connection of $V(q)$ may have different lengths on the boundaries of two cylinders or minimal domains, it is necessary to modify the leaves joining singularities such that the adjacent cylinders or minimal domains match along the leaves. This can be achieved by introducing additional singularities and a saddle connection. Suppose that a saddle connection has length $\ell_{i}$ on one boundary and $\ell_{j}$ on the other boundary, with $\ell_{i}>\ell_{j}$. Consider the saddle connection of length $\ell_{i}$. We pick three points on this saddle connection such that it is divided into four arcs, with the two middle subarcs having equal length $\frac{\ell_{i}-\ell_{j}}{2}$. By identifying these two middle subarcs, we obtain a saddle connection joining a one-prong singularity to a three-prong singularity (see Figure 1). Since the remaining length of the saddle connection is $\ell_{j}$, the two corresponding components of $V(q)$ can be glued together along their boundaries.
$\ell_{i}$
$\ell_{j}$
Figure 1. After rescaling the leaves, a saddle connection may have two different lengths on the boundaries of two adjacent components of $V(q)$. These components are glued together by introducing an additional saddle connection (right).
After rescaling each leaf of $V(q)$ by the corresponding factor, we glue together all cylinders and minimal domains of $V(q)$ along their boundaries, introducing saddle connections when necessary. In this way, we obtain a measured foliation $V^{\prime}$ from $V(q)$. It is clear that $V^{\prime}$ is Whitehead equivalent to $V(q)$, since $V^{\prime}$ is obtained from $V(q)$ by a sequence of Whitehead moves. In addition, we construct a measured foliation $F^{\prime}$ by gluing the leaves of $H(q)$, restricted to the cylinders and minimal domains of $V(q)$, along their boundaries. The transverse measure of $F^{\prime}$ is determined by the lengths of arcs on the leaves of $V^{\prime}$. Consequently, we obtain a measured foliation $F^{\prime}$ from $H(q)$ satisfying that
$a_{j}i(G_{j},F^{\prime})=c_{j}i(G_{j},H(q))$
for each $j=1,\cdots,n$. Then,
@@ 4. The asymptotic behavior of pairs of Teichmüller rays
If two measured foliations $F,F^{\prime}\in\mathcal{MF}$ are absolutely continuous, it is evident that
@@
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray determined by a unit-norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. We define a function $\mathcal{E}_{q}:\mathcal{MF}\to\mathbb{R}_{\geq 0}$ associated with the Teichmüller ray, which is given by
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit-norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
Lemma 4.2 ([21, Lemma 29]).
Let $q$ and $q^{\prime}$ be two unit-norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}>0$. If the vertical foliation of $q^{\prime}$ can be expressed as $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$, where $b_{j}\geq 0$, then
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit-norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
Let $q$ and $q^{\prime}$ be two unit-norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
@@
where $\mathcal{MF}^{\prime}_{0}(V(q))=\{F\in\mathcal{MF}\mid F=\sum_{j=1}^{n}c_{j}G_{j},\;c_{j}>0\}$.
@@
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit-norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then